# arXiv:2603.02406v1[cs.LG] 02 Mar 2026

### 02 Mar 2026 v1

@@ 5 Discussion
RigidSSL-Perturb Improves Geometric and Physical Quality of Learned Structures. In the self-supervised pretraining step in Phase I, we employ data augmentation by adding noise to the original geometry to construct a perturbed geometry as the second view. This choice is motivated by three considerations. (1) Gaussian perturbation can be interpreted as a form of “masking” applied to atom positions in 3D Euclidean space. Analogous to how self-supervised learning in computer vision masks image patches or in NLP masks tokens, noise injection in geometry partially obscures atomic coordinates and encourages the model to infer consistent global features across corrupted inputs. (2) At small noise scales, this augmentation mimics systematic errors that arise during data collection or prediction. Protein structures from crystallography, cryo-EM, or computational databases such as AFDB often contain small uncertainties in atom placement. Exposing the model to controlled perturbations helps it become robust to such imperfections. (3) Gaussian perturbations can also approximate natural conformational fluctuations around a stable equilibrium. Proteins are dynamic molecules, and their functional states often involve small deviations from the ground-state structure. Training with noisy perturbations introduces the model to this variability, allowing it to capture flexible yet physically plausible structural patterns.
@@
As shown in Table˜1, Phase I pretraining improves designability, indicating that the model’s outputs shift toward a more reliable distribution, but this comes at the cost of reduced diversity. While random perturbation provides clear benefits for robustness and representation learning, it also introduces an inherent trade-off between designability and diversity. By enforcing invariance across noisy views, the model learns to emphasize stable, fold-defining features while disregarding minor structural deviations. This bias enhances designability, as downstream generators built on such representations tend to produce more foldable backbones. However, the same invariance can also suppress conformational variability, collapsing nearby structures into a single representation and thereby reducing diversity. In practice, noise perturbation prioritizes stability over exploration: it improves the reliability of designed proteins but may limit the diversity of generated folds. In Appendix˜G, we analyze how the noise scales in RigidSSL-Perturb affect the structures generated in downstream tasks.
@@
RigidSSL-MD Improves Conformational Flexibility and Structural Diversity. In the second phase, RigidSSL-MD incorporates MD trajectories to build augmented views. While this improves structural diversity, it reduces designability compared to perturbation-based pretraining (Table˜1). The trade-off arises because MD pretraining is not purely self-supervised: it depends on force-field energies and gradients, inheriting biases that may misalign with de novo design goals and even cause negative transfer, as seen in prior studies (Liu et al., 2023a). At the same time, MD trajectories capture conformational dynamics across multiple metastable states, enriching structural flexibility and promoting greater diversity in the learned representations. However, because most structure prediction models are trained on stable, static conformations, incorporating such dynamic biases can sometimes hinder re-folding performance. Overall, MD-based pretraining emphasizes flexibility and diversity at the expense of foldability, whereas perturbation-based SSL favors stability and designability. These two paradigms may therefore serve as complementary strategies, depending on the downstream design objectives.

### 06 Mar 2026 v2

# Rigidity-Aware Geometric Pretraining for Protein Design and Conformational Ensembles

@@ 5 Discussion
RigidSSL-Perturb Improves Geometric Quality of Learned Structures. In the self-supervised pretraining step in Phase I, we employ data augmentation by adding noise to the original geometry to construct a perturbed geometry as the second view. This choice is motivated by three considerations. (1) Gaussian perturbation can be interpreted as a form of “masking” applied to atom positions in 3D Euclidean space. Analogous to how self-supervised learning masks image patches in computer vision or tokens in NLP, noise injection in geometry partially obscures atomic coordinates and encourages the model to infer consistent global features across corrupted inputs. (2) At small noise scales, this augmentation mimics systematic errors that arise during data collection or prediction. Protein structures from crystallography, cryo-EM, or computational databases such as AFDB often contain small uncertainties in atom placement. Exposing the model to controlled perturbations helps it become robust to such imperfections. (3) Gaussian perturbations can also approximate natural conformational fluctuations around a stable equilibrium. Proteins are dynamic molecules, and their functional states often involve small deviations from the ground-state structure. Training with noisy perturbations introduces the model to this variability, allowing it to capture flexible yet physically plausible structural patterns.
@@
As shown in Table˜1, Phase I pretraining improves designability, indicating that the model’s outputs shift toward a more reliable distribution, but this comes at the cost of reduced diversity compared to Phase II. By maximizing mutual information between noisy views, the model learns representations that emphasize stable, fold-defining features while downweighting minor structural deviations. Downstream generators built on such representations therefore tend to produce more reliable and foldable backbones. However, this objective can also suppress conformational variability by collapsing nearby structures into a single representation, thereby limiting the diversity of generated folds.
@@
RigidSSL-MD Promotes Structural Diversity and Biophysical Fidelity. In the second phase, RigidSSL-MD incorporates MD trajectories to build augmented views. While this improves structural diversity, it reduces designability compared to perturbation-based pretraining (Table˜1). The trade-off arises because MD pretraining is not purely self-supervised in the sense that training views are drawn from force-field-generated trajectories, inheriting biases that may misalign with de novo design goals and even cause negative transfer, as seen in prior studies (Liu et al., 2023a). In contrast, the same physically grounded bias improves higher-order biophysical properties (Table˜6). Another explanation for the lower designability is that most structure prediction models are trained on stable, static conformations. While MD trajectories capture conformational dynamics to enrich structural flexibility, they also bias the model toward generating metastable conformations rather than the absolute ground states expected by folding and inverse-folding pipelines, thereby hindering re-folding performance. Overall, RigidSSL-MD emphasizes diversity and physical fidelity, whereas RigidSSL-Perturb favors geometric quality described by the designability metric. These two paradigms may therefore serve as complementary strategies, depending on the downstream design objectives.